# A decade of the fast-varying ionospheric and magnetospheric magnetic fields

### v1

@@
A decade of the fast-varying ionospheric and magnetospheric magnetic fields
@@ 2 Institute of Geophysics and Meteorology
Ionosphere, Magnetosphere, Geomagnetism, Space weather
@@
The time-varying geomagnetic field is a superposition of contributions from multiple internal and external current systems. A major source of geomagnetic variations at periods less than few years are current systems external to the solid Earth, namely the ionospheric and magnetospheric currents, as well as associated mantle-induced currents. The separation of these three sources is mathematically underdetermined using either ground or satellite measurements alone, but becomes tractable when the two datasets are combined. Based on this concept, we developed a new geomagnetic field modelling approach that allows us to simultaneously characterise the mid-latitude ionospheric, magnetospheric and the internal induced magnetic fields using ground and satellite observations for all local times and magnetic conditions, and without prescribing any harmonic behaviour on these current systems in time, as is typical in other models. By applying this new method to a 10-year dataset of ground observatory and multi-satellite measurements from 2014 to 2023, we obtained the time series of the spherical harmonic coefficients of the ionospheric, magnetospheric and induced fields. These new time series unravel complex non-periodic dynamics of the external magnetic fields during global geomagnetic storms, as well as periodicities in the magnetospheric coefficients linked to solar activities and periodic ionospheric magnetic fields linked to lunar daily variations, contributing to a better picture of the dynamics of the external currents and magnetosphere-ionosphere interactions. Finally, new approach provides more consistent field estimates for electromagnetic induction studies, and, owing to its flexibility and on-the-fly updates, can facilitate space weather operations.
@@ 1 Introduction
Time variations of magnetic field observed on the ground or at a spacecraft represent a superposition of contributions from multiple time-varying current systems. The electric current generated by the dynamo action in the fluid outer core is responsible for sustaining the main field over geological time scales, and is also the major contributor to magnetic field variations at long periods, known as secular variations (Backus et al., 1996). Meanwhile, the extraneous current systems in the ionosphere and the magnetosphere generate time-varying magnetic field, which then induce currents in the electrically conductive solid Earth (Schuster, 1889; Olsen, 1999a). Characterising the ionospheric and magnetospheric contributions to the observed magnetic field is of interest for several reasons. First, magnetic fields of ionospheric and magnetospheric origins are used to quantitatively study respective electric current systems, and serve as a proxy for understanding the underlying dynamics in these regions (Yamazaki & Maute, 2017; Tsyganenko, 2019; Laundal et al., 2021). On the other hand, the relation between these magnetic fields of external origin and their induced counterparts in the Earth interior can be used to probe the electrical conductivity in the Earth’s subsurface (Olsen, 1999b; Kelbert et al., 2009; Kuvshinov, 2012; Grayver, 2024). More accurate space physics and subsurface conductivity models are then essential for a practically useful space weather hazard evaluation, as the latter is driven by the dynamics in the extraneous current systems (Tóth et al., 2005; Pulkkinen et al., 2013) but sensitive to the subsurface electrical conductivity as well (Juusola et al., 2020; Kelbert & Lucas, 2020). Last but not least, better characterisation of the ionospheric and magnetospheric systems also improves the core field modelling (Finlay et al., 2017), and indirectly leads to improvements in crust and ocean field models (Olsen et al., 2017; Grayver et al., 2024).
@@
Noting the problems and limitations with geomagnetic ground observatories, recent studies increasingly incorporate satellite measurements in geomagnetic field modelling. Continuous magnetic field observations from Low-Earth-Orbit (LEO) satellites have enabled significant advances in geomagnetic field modelling, leading to improved spatial resolution of various magnetic field models, including for the slow-varying or static core (Alken et al., 2021), crustal (Maus et al., 2007; Olsen et al., 2017) and oceanic (Sabaka et al., 2020; Grayver et al., 2024) fields. This also holds for models of ionospheric and magnetospheric fields (Lesur et al., 2010; Chulliat et al., 2016; Sabaka et al., 2018; Finlay et al., 2020; Baerenzung et al., 2020). However, the second problem regarding separation of external field sources is still mostly at large. Most external field models attempt to achieve this by reducing the spatial or/and temporal complexity of the external field model parametrisations. For instance, by taking only the dark-side magnetic observations, these models attribute the magnetic field external to observatories and satellites to the magnetosphere, assuming ionospheric magnetic fields can be neglected at night (Lesur et al., 2010; Finlay et al., 2020; Baerenzung et al., 2020). This approach limits the usable data to only the dark-side field measurements, typically discarding more than half of the available data. In addition, even the dark-side data may contain ionospheric contributions. Apart from the potential dynamics in the F region during the night, the transient behaviour of electromagnetic induction also means that the dark-side data contain mantle signals that are induced by the day-time ionosphere and persist into the night (Maus & Weidelt, 2004; Grayver et al., 2021). Another common practice is to use only observations taken during quiet magnetic conditions. For instance, a dedicated magnetic field model of the ionosphere and equatorial electrojet (EEJ) by Chulliat et al. (2016) delivers a model of the mid-latitude ionosphere during quiet phases. Although the model has a high spatial resolution, the temporal dynamics of the ionosphere is approximated by a small set of daily and seasonal time harmonics. Other temporal components are approximated by modulating the predicted field with an external time-varying proxy (e.g. F10.7 flux). More elaborate Comprehensive Inversion (CI) approach tries to parametrise the ionospheric magnetic field together with other components, including a lowest-degree magnetospheric field (Sabaka et al., 2018). While the parametrisation of ionospheric currents in CI is quite elaborate and fits observations at mid and equatorial latitudes well, a small set of discrete points in the frequency domain at annual, semiannual and diurnal bands also means that other temporal variations are not captured. Similarly, the model is constructed using only quiet magnetic periods. Further, these models are difficult to update on the fly as small volumes of new observations are added in time, limiting their applications for operational space-weather or for studying individual events, such as geomagnetic storms.
@@
To overcome the limitations of current models, and to better exploit the increasing number of satellite observations delivered by dedicated missions and platform magnetometers, we present here a novel approach to simultaneously characterise the ionospheric, magnetospheric and mantle-induced magnetic fields using all available ground and space observations and high temporal resolution. Our approach utilises the geometrical configuration of ground and low Earth orbit (LEO) geomagnetic observations. Indeed, the ground observatories alone cannot distinguish between the ionospheric and magnetospheric contributions, as both are external to the observations. The introduction of satellite observations in the region between the ionosphere and magnetospheric currents, however, renders the problem feasible. We construct model in short time-bins, allowing for a continuous Swarm-era model, including all local times and magnetic activity conditions. In addition to ionospheric and magnetospheric components, we estimate time series of SH coefficients for the field induced within the solid Earth, rendering the model output also relevant for mantle induction studies.
@@ 2 Methodology
The variation of the magnetic field at location $\mathbf{r}$ and time $t$ is described by the Maxwell’s equations, which read
@@
The SH representation has long been used for constructing geomagnetic field models from ground-based observations (Gauss, 1877). Since the ground observations are made in the electrically insulating neutral atmosphere, the potential formulation of the magnetic field in a spherical shell is valid. In recent decades, an increasing number of studies also included satellite observations in the same SH representation (Bloxham & Jackson, 1989; Langlais & Mandea, 2000; Olsen et al., 2006; Baerenzung et al., 2020). This approach is justified since satellite observations taken at $400-800$ km altitude are away from major source regions: the magnetospheric current systems (Ganushkina et al., 2018), and the E-region dynamo in the ionosphere (Yamazaki & Maute, 2017).
@@
In this study, we are interested in the magnetic fields that are ultimately attributed to extraneous current systems. To this purpose, we consider three major electrical current systems contributing to the magnetic field: 1) the magnetospheric currents, most dominant of which is the ring current, 2) the ionospheric currents, e.g. from the E-region dynamo, and 3) their induced counterparts in the solid Earth. As we are most interested in the large-scale features in the magnetic field, we shall restrict the ionospheric currents to the mid-latitude currents. This excludes the equatorial and polar current systems, which require dedicated parametrisations. Meanwhile, the core field with its secular variations and the lithospheric field are considered removed from the system. The geometric setup of our model is illustrated in Fig. 1.
@@
Under these considerations, the observatory Gauss coefficients of internal origin represent contributions from electric currents induced in the Earth’s interior by nature of electromagnetic induction. We write these coefficients alternatively as
@@
Using $B_{r}=-\partial_{r}V_{\mathrm{obs}}$ for $r<a+h$ and $B_{r}=-\partial_{r}V_{\mathrm{sat}}$ for $r>a+h$, a one-to-one relation between the ionospheric Gauss coefficients can be established
@@
Combining eq.(12) with eqs.(7-10), we can rewrite the SH representation for the ground observations (5) as
@@
These representations reduce the magnetic potential to the source-specific Gauss coefficients $(g_{nm}^{\mathrm{int}},h_{nm}^{\mathrm{int}})$, $(q_{nm}^{\mathrm{ion}},s_{nm}^{\mathrm{ion}})$, $(q_{nm}^{\mathrm{mag}},s_{nm}^{\mathrm{mag}})$, hereinafter referred to as mantle induced coefficients, ionospheric coefficients and magnetospheric coefficients, respectively. These Gauss coefficients can be estimated by minimising the data misfit of the combined satellite and ground observatory dataset. The compilation and pre-processing of the dataset and the model estimation process are detailed in Sections (3) and (4), respectively.
@@ 3 Data
While Eq. (13-14) provides magnetic field models that are continuous in both space and time, we implement a time-discrete version by estimating the Gauss coefficients within non-overlapping time bins. For this purpose, the ground and satellite observations must be both arranged into separate time bins, within which the corresponding SH coefficients are assumed to be constant. This inevitably introduces a trade-off between spatial coverage and the temporal resolution of the satellite data. We found that a time bin of $3$ hours provides a fair balance between the spatial and temporal resolution. The model parametrisation can be easily adjusted depending on future applications and availability of new satellite missions and observatories.
@@ 3.1 Satellite data
We then filter out satellite observations taken at locations with an absolute quasi-dipole (QD) latitude (Richmond, 1995) below $5^{\circ}$ or above $56^{\circ}$. This is to avoid the effects of equatorial and polar electrojets, which are localised current systems requiring dedicated parametrisations and cannot be well described by low-degree spherical harmonics. Contrary to approaches deriving full geomagnetic models with a focus on core models, such as the CHAOS model (Finlay et al., 2020) or the Comprehensive Model (Sabaka et al., 2020), we do not filter observations based on solar zenith angle or magnetic indices, nor do we parametrise the ionospheric field using specific diurnal/seasonal time harmonics. Since the spacial and temporal behaviour of the external field is of prime interest here, we impose no constraint on the magnetic conditions and inject no prior knowledge on the periodicity of reconstructed sources. Instead, the parameter estimation is carried out independently for each time window, and the temporal behaviour will be analysed in a completely a posteriori fashion. A significant advantage of such model is the ease of its update, which can be performed on the fly in real-time (limited only by the geomagnetic data latency).
@@ 3.2 Ground observatory data
We compiled the vector magnetic field data measured by ground observatories from the INTERMAGNET and World Data Centre for Geomagnetism (WDC) within the same time period (INTERMAGNET; et al., 2024). The hourly mean time series are binned in $3$-hr time bins for consistency with the satellite data. The same filtering based on QD latitude ($5^{\circ}-56^{\circ}$ north and south) has been applied. The distribution of ground observatories used in this study is shown in Fig. (2).
@@ 3.2.1 Observatory bias
Since the observatory bias cannot be co-estimated in our model with a piece-wise constant temporal parametrisation of the ionosphere, we derive the observatory biases in a separate step. We took ground observatory and satellite data from magnetic quiet times (defined as times when $K_{p}\leq 2$ and $|dDst/dt|\leq 3\mathrm{nT/h}$) with the Sun at least $10^{\circ}$ below the horizon. Data fulfilling these conditions are referred to as quiet night data. The purpose of choosing only quiet night data is to minimise the effects of the ionosphere. The quiet night data are then fitted using a magnetic field model with only magnetospheric field and the mantle induced field, and the observatory bias consisting of one time independent vector per observatory. Additional vector biases are introduced for observatories where large offsets are seen in different segments of observation. Due to the limited spatial coverage of the filtered data, the magnetospheric field and the mantle induced field are parametrised only up to SH degree and order one. The bias values are then determined using a computation- and memory-efficient linear least squares algorithm for block-diagonal matrix with intercept (Appendix C).
The estimated observatory biases are shown in supplementary Fig. LABEL:fig:obs-bias-est-map. The observatory biases derived in this study agree well with the CI-derived biases, and both are close to the quiet night mean values (Fig. LABEL:fig:obs-bias-est, LABEL:fig:obs-bias-est-diffmean). The reason to carry out this estimation instead of using the CI-derived biases is to retain the consistency with our data set, and with our model which does not prescribe a periodic behaviour of the ionospheric field, whereas CI does. These vector biases are then removed from the ground observatory measurements.
@@
Figure 3: Average number of data points per time window as a function of time. Each vector component is treated as an independent datum.

### v2

@@
A decade of the fast-varying ionospheric and magnetospheric magnetic fields from ground and multi-satellite observations
@@ 2 Institute of Geophysics and Meteorology
Geomagnetic induction; Magnetic field variations through time; Satellite magnetics; Fourier analysis
@@
The time-varying geomagnetic field is a superposition of contributions from multiple internal and external current systems. A major source of geomagnetic variations at periods less than a few years are current systems external to the solid Earth, namely the ionospheric and magnetospheric currents, as well as associated induced currents. The separation of these three sources is mathematically underdetermined using either ground or satellite measurements alone, but becomes tractable when the two datasets are combined. Based on this concept, we developed a new geomagnetic field modelling approach that allows us to simultaneously characterise the mid-latitude ionospheric, magnetospheric and the internal induced magnetic fields using ground and satellite observations for all local times and magnetic conditions, and without prescribing any harmonic behaviour on these current systems in time, as is typical in other models. By applying this new method to a 10-year dataset of ground observatory and multi-satellite measurements from 2014 to 2023, we obtained the time series of the spherical harmonic coefficients of the ionospheric, magnetospheric and induced fields. These new time series allow the study of complex non-periodic dynamics of the external magnetic fields during global geomagnetic storms, as well as periodicities in the magnetospheric coefficients linked to solar activities and periodic ionospheric magnetic fields linked to lunar daily variations, contributing to a more complete picture of the dynamics of the external currents and magnetosphere-ionosphere interactions, and facilitating more accurate space weather nowcast and forecast. Finally, the new approach allows for a better characterisation of internal induced field sources, leading to higher quality electromagnetic transfer functions.
@@ 1 Introduction
Time variations of the magnetic field observed on the ground or at a spacecraft represent a superposition of contributions from multiple time-varying current systems. The electric current generated by the dynamo action in the fluid outer core is responsible for sustaining the main field over geological time scales, and is also the major contributor to magnetic field variations at long periods, known as secular variations (Backus et al., 1996). Meanwhile, the extraneous current systems in the ionosphere and the magnetosphere generate time-varying magnetic fields, which then induce currents in the electrically conductive solid Earth (Schuster, 1889; Olsen, 1999a). Characterising the ionospheric and magnetospheric contributions to the observed magnetic field is of interest for several reasons. First, magnetic fields of ionospheric and magnetospheric origins are used to quantitatively study respective electric current systems, and serve as a proxy for understanding the underlying dynamics in these regions (Yamazaki & Maute, 2017; Tsyganenko, 2019; Laundal et al., 2021). On the other hand, the relation between these magnetic fields of external origin and their induced counterparts in the Earth interior can be used to probe the electrical conductivity in the Earth’s subsurface (Olsen, 1999b; Kelbert et al., 2009; Kuvshinov, 2012; Grayver, 2024). More accurate space physics and subsurface conductivity models are then essential for a practically useful space weather hazard evaluation, as the latter is driven by the dynamics in the extraneous current systems (Tóth et al., 2005; Pulkkinen et al., 2013) but sensitive to the subsurface electrical conductivity as well (Juusola et al., 2020; Kelbert & Lucas, 2020). Last but not least, better characterisation of the ionospheric and magnetospheric systems also improves the core field modelling (Finlay et al., 2017), and indirectly leads to improvements in crust and ocean field models (Olsen et al., 2017; Grayver et al., 2024).
@@
Noting the problems and limitations with geomagnetic ground observatories, recent studies increasingly incorporate satellite measurements in geomagnetic field modelling. Continuous magnetic field observations from Low-Earth-Orbit (LEO) satellites have enabled significant advances in geomagnetic field modelling, leading to improved spatial resolution of various magnetic field models, including those for the slow-varying or static core (Alken et al., 2021), crustal (Maus et al., 2007; Olsen et al., 2017) and oceanic (Sabaka et al., 2020; Grayver et al., 2024) fields. This also holds for models of ionospheric and magnetospheric fields (Lesur et al., 2010; Chulliat et al., 2016; Sabaka et al., 2018; Finlay et al., 2020; Baerenzung et al., 2020). However, the second problem regarding separation of external field sources is still mostly at large. Most external field models attempt to achieve this by reducing the spatial and/or temporal complexity of the external field model parametrisations. For instance, by taking only the dark-side magnetic observations, these models attribute the magnetic field external to observatories and satellites to the magnetosphere, assuming ionospheric magnetic fields can be neglected at night (Lesur et al., 2010; Finlay et al., 2020; Baerenzung et al., 2020). This approach limits the usable data to only the dark-side field measurements, typically discarding more than half of the available data. In addition, even the dark-side data may contain ionospheric contributions. Apart from the potential dynamics in the F region during the night, the transient behaviour of electromagnetic induction also means that the dark-side data contain internal signals that are induced by the day-time ionosphere and persist into the night (Maus & Weidelt, 2004; Grayver et al., 2021). Another common practice is to use only observations taken during quiet magnetic conditions. For instance, a dedicated magnetic field model of the ionosphere and equatorial electrojet (EEJ) by Chulliat et al. (2016) delivers a model of the mid-latitude ionosphere during quiet phases. Although the model has a high spatial resolution, the temporal dynamics of the ionosphere is approximated by a small set of daily and seasonal time harmonics. Other temporal components are approximated by modulating the predicted field with an external time-varying proxy (e.g. F10.7 flux). More elaborate Comprehensive Inversion (CI) approach tries to parametrise the ionospheric magnetic field together with other components, including a lowest-degree magnetospheric field (Sabaka et al., 2018). While the parametrisation of ionospheric currents in CI is quite elaborate and fits observations at mid and equatorial latitudes well, a small set of discrete points in the frequency domain at annual, semiannual and diurnal bands also means that other temporal variations are not captured. Similarly, the model is constructed using only quiet magnetic periods. Further, these models are difficult to update on the fly as small volumes of new observations are added in time, limiting their applications for operational space-weather or for studying individual events, such as geomagnetic storms.
@@
To overcome the limitations of current models, and to better exploit the increasing number of satellite observations delivered by dedicated missions and platform magnetometers, we present here a novel approach to simultaneously characterise the ionospheric, magnetospheric and internally induced magnetic fields using all available ground and space observations and high temporal resolution. Our approach utilises the geometrical configuration of ground and low Earth orbit (LEO) geomagnetic observations. Indeed, the ground observatories alone cannot distinguish between the ionospheric and magnetospheric contributions, as both are external to the observations. The introduction of satellite observations in the region between the ionosphere and magnetospheric currents, however, renders the problem feasible. We construct model in short time-bins, allowing for a continuous Swarm-era model, including all local times and magnetic activity conditions. In addition to ionospheric and magnetospheric components, we also estimate time series of the spherical harmonic coefficients of the internally induced field, rendering the model output also relevant for mantle induction studies.
@@ 2 Methodology
The time-dependent magnetic field at location $\mathbf{r}$ and time $t$ is described by the Maxwell’s equations, which read
@@
The SH representation has long been used for constructing geomagnetic field models from ground-based observations (Gauss, 1839). Since the ground observations are made in the electrically insulating neutral atmosphere, the potential formulation of the magnetic field in a spherical shell is valid. In recent decades, an increasing number of studies also included satellite observations in the same SH representation (Langel et al., 1980; Bloxham & Jackson, 1989; Langlais & Mandea, 2000; Olsen et al., 2006; Baerenzung et al., 2020). This approach is justified since satellite observations taken at $400-800$ km altitude are away from major source regions: the magnetospheric current systems (Ganushkina et al., 2018), and the E-region dynamo in the ionosphere (Yamazaki & Maute, 2017).
@@
In this study, we are interested in the magnetic fields that are ultimately attributed to extraneous current systems. For this purpose, we consider three major electrical current systems contributing to the magnetic field: 1) the magnetospheric currents, the most dominant of which is the ring current, 2) the ionospheric currents, mostly from the E-region dynamo, and 3) the corresponding induced signals originating in the solid Earth or in the ocean, and hence internal to all observations. As we are most interested in the large-scale features in the magnetic field, we shall restrict the ionospheric currents to the mid-latitude currents. This excludes the equatorial and polar current systems, which require dedicated parametrisations. Meanwhile, the core field with its secular variations and the lithospheric field are considered removed from the system. The geometric setup of our model is illustrated in Fig. 1.
@@
Under these considerations, the observatory Gauss coefficients of internal origin represent contributions from electric currents induced in the Earth’s interior or in the ocean by nature of electromagnetic induction. We write these coefficients alternatively as
@@
Using $B_{r}=-\partial_{r}V_{\mathrm{obs}}$ for $r<a+h$ and $B_{r}=-\partial_{r}V_{\mathrm{sat}}$ for $r>a+h$, a one-to-one relation between the ionospheric Gauss coefficients can be established (see also Sabaka et al., 2002)
@@
Combining Eqs. (7-10) with (12), we can rewrite the SH representation for ground observations (5) as
@@
These representations reduce the magnetic potential to the source-specific Gauss coefficients $(g_{nm}^{\mathrm{int}},h_{nm}^{\mathrm{int}})$, $(q_{nm}^{\mathrm{ion}},s_{nm}^{\mathrm{ion}})$, $(q_{nm}^{\mathrm{mag}},s_{nm}^{\mathrm{mag}})$, hereinafter referred to as internally induced coefficients, ionospheric coefficients and magnetospheric coefficients, respectively. These Gauss coefficients can be estimated by minimising the data misfit of the combined satellite and ground observatory dataset. The compilation and pre-processing of the dataset and the model estimation process are detailed in Sections 3 and 4, respectively.
@@ 3 Data
While Eqs. (13-14) provides magnetic field models that are continuous in both space and time, we implement a time-discrete version by estimating the Gauss coefficients within non-overlapping time bins. For this purpose, the ground and satellite observations must be both arranged into separate time bins, within which the corresponding SH coefficients are assumed to be constant. This inevitably introduces a trade-off between spatial coverage and the temporal resolution of the satellite data. We found that a time bin of $3$ hours provides a fair balance between the spatial and temporal resolution. The model parametrisation can be easily adjusted depending on future applications and availability of new satellite missions and observatories.
@@ 3.1 Satellite data
We then filter out satellite observations taken at locations with an absolute quasi-dipole (QD) latitude (Richmond, 1995) below $5^{\circ}$ or above $56^{\circ}$. This is to avoid the effects of equatorial and polar electrojets, which are localised current systems requiring dedicated parametrisations and cannot be well described by low-degree spherical harmonics. Contrary to approaches deriving full geomagnetic models with a focus on core models, such as the CHAOS model (Finlay et al., 2020) or the Comprehensive Model (Sabaka et al., 2020), we do not filter observations based on solar zenith angle or magnetic indices, nor do we parametrise the ionospheric field using specific diurnal/seasonal time harmonics. Since the spatial and temporal behaviour of the external field is of prime interest here, we impose no constraint on the magnetic conditions and inject no prior knowledge on the periodicity of reconstructed sources. Instead, the parameter estimation is carried out independently for each time window, and the temporal behaviour will be analysed in a completely a posteriori fashion. A significant advantage of such model is the ease of its update, which can be performed on the fly in real-time (limited only by the geomagnetic data latency).
@@ 3.2 Ground observatory data
We compiled the vector magnetic field data measured by ground observatories from the INTERMAGNET and World Data Centre for Geomagnetism (WDC) within the same time period (INTERMAGNET; et al., 2024). The hourly mean time series are binned in $3$-hr time bins for consistency with the satellite data. The same filtering based on QD latitude ($5^{\circ}-56^{\circ}$ north and south) has been applied. The distribution of ground observatories used in this study is shown in Fig. 2.
@@ 3.2.1 Observatory bias
Since the observatory bias cannot be co-estimated in our model with a piece-wise constant temporal parametrisation of the ionosphere, we derive these biases in a separate step. We took ground observatory and satellite data from magnetic quiet times (defined as times when $K_{p}\leq 2$ and $|dDst/dt|\leq 3\mathrm{nT/h}$) with the Sun at least $10^{\circ}$ below the horizon. Data fulfilling these conditions are referred to as quiet night data. The quiet night data are then fitted using a magnetic field model with only magnetospheric field and the internally induced field, and the observatory bias consisting of one time independent vector per observatory. Additional vector biases are introduced for observatories where large offsets are seen in different segments of observation. Due to the limited spatial coverage of the filtered data, the magnetospheric field and the internally induced field are parametrised only up to SH degree and order one. Although the dark-side data are not fully free of ionospheric and its corresponding induced field contributions, such effects are minimal due to much weaker ionospheric currents (Price & Wilkins, 1963). Using a parametrisation without the ionosphere on the quiet night data is a necessary compromise since the addition of the ionospheric source would lead to an intrinsic non-uniqueness in the bias estimation problem (Appendix B). In addition, note that the nullspace of the forward problem with the ionosphere is characterised by a static field. Therefore, the omission of the ionosphere in the bias estimation can only introduce a small static offset for the entire data span (10 years in our model), which does not interfere with the temporal variability of the field estimates.
The bias values are then determined using a computation- and memory-efficient linear least squares algorithm for block-diagonal matrix with intercept (Appendix C). The estimated observatory biases are shown in supplementary Fig. LABEL:fig:obs-bias-est-map. The observatory biases derived in this study agree well with the CI-derived biases, and both are close to the quiet night mean values (Figs. LABEL:fig:obs-bias-est, LABEL:fig:obs-bias-est-diffmean, Table LABEL:tab:bias-rmsd). The reason to carry out this estimation instead of using the CI-derived biases is to retain the consistency with our data set, and with our model which does not prescribe a periodic behaviour of the ionospheric field, whereas CI does. These vector biases are then removed from the ground observatory measurements.
@@
Figure 3: Average number of data points per time bin as a function of time. Each vector component is treated as an independent datum.